# Relaying Quantum Information

### v1

@@ IV Main Results
Here we present our main results for the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We establish achievable rates that are based on different coding strategies, with and without the entanglement assistance.
@@ IV-A Rate-Limited Entanglement Assistance
In the decode-forward case, the relay decodes the entire information. Consider the three party communication scheme shown in Fig. 3. Let $\mathcal{N}_{AD\rightarrow BE}$ be a quantum channel, and let ${|{\sigma}\rangle}_{AA_{1}D}$ be any pure state with $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$.
@@
Assume that Bob has access to entanglement assistance represented by entanglement consumption and generation rates $L_{B}$ and $\hat{L}_{B}$, respectively. Then any rate point $(Q,L_{B},\hat{L}_{B})$ such that:
$\begin{split}Q\leq H(A_{1}|D)_{\sigma}-L_{B}+\hat{L}_{B}\\ Q\leq I(A_{1}\rangle E)_{\omega}-L_{B}-\hat{L}_{B}\\ Q\leq I(A_{1}\rangle B)_{\omega}+L_{B}-\hat{L}_{B},\end{split}$ (8)
is achievable for transmission with rate-limited entanglement assistance, where $A_{1}$ is an arbitrary auxiliary system.
The first condition states that the quantum information transmitted by Alice, along with the entanglement consumed from her share of the EPR pairs, must fit within the channel input available to Alice, conditioned on the relay’s input to the channel. Specifically, let $\Delta L_{B}=L_{B}-\hat{L}_{B}$ denote the net entanglement consumption. Then, the first condition states that $Q+\Delta L_{B}\leq H(A_{1}|D)_{\sigma}$. The third condition says that the channel can transmit $Q-\Delta L_{B}\leq I(A_{1}\rangle B)$ qubits per channel use, when taking the cost of entanglement into account, hence subtracting the net entanglement consumption. The second condition directly reflects the bottleneck introduced by the relay, where both $L_{B}$ and $\hat{L}_{B}$ appear with the negative sign. This arises due to the nature of our Block Markov coding scheme, where the entanglement assistance available to Bob in block $j$ is effectively generated by the relay in block $j-1$. See Sec. VI-B.
IV-B Without Entanglement Assistance
Our main result is given below. Define
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\min\left\{I(A_{1}\rangle B)_{\omega},I(A_{1}\rangle E)_{\omega}\right\}\right],$
where the maximum is taken over all pure states of the form ${|{\sigma_{A_{1}AD}}\rangle}$, and where $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$. The subscript ‘D-F’ stands for decode-forward bound.
@@ IV-B Without Entanglement Assistance
$C(\mathcal{N})\geq\mathsf{Q}_{\text{D-F}}(\mathcal{N}).$ (9)
IV-C Unlimited Entanglement Assistance
In the next theorem, we consider communication of quantum information via the quantum relay channel, when Alice and Bob are provided with pre-shared entanglement assistance, in which case, $L_{B}>\hat{L}_{B}>0$ (see Proposition 1). Let
$\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\frac{1}{2}I(A_{1};B)_{\omega}-\frac{1}{2}I(A_{1};D)_{\sigma}\right],$
where the maximization is as before, and the subscript ‘EA’ indicates the availability of entanglement assistance.
Theorem 3.
$C_{\text{EA}}(\mathcal{N})\geq\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N}).$ (10)
@@ IV-C Unlimited Entanglement Assistance
The theorem above has an immediate consequence on the transmission of classical information with entanglement assistance. Specifically, using the superdense coding protocol, classical information can be sent at a rate of $R\geq$ $2\cdot\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})$ (classical information bits per transmission).
@@
The input state ${|{\sigma}\rangle}_{AA_{1}D}$ represents the correlation that is generated between the transmitter and the relay. Intuitively, we think of the auxiliary system $A_{1}$ as Alice’s entanglement resource. Due to entanglement monogamy, if the resource $A_{1}$ is maximally entangled with the relay’s transmission, $D$, then there is no correlation with the channel input, $A$. Similarly, if the resource is maximally entangled with the channel input, then it has no correlation with the relay. Therefore, it is not immediately obvious how to distribute the correlation between $A_{1}$, $A$, and $D$.
Taking ${|{\sigma}\rangle}_{AA_{1}D}={|{\phi_{A_{1}A}}\rangle}\otimes{|{\psi}\rangle}_{D}$ yields the direct transmission bound:
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\geq\max_{\sigma}\left[I(A_{1}\rangle B)_{\omega}\right]$
i.e., the coherent information of the channel from Alice to Bob. Similarly, we recover the entanglement-assisted-capacity:
$\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})\geq\max_{\sigma}\left[I(A_{1};B)_{\omega}\right].$
On the other hand, ${|{\sigma}\rangle}_{AA_{1}D}={|{\psi}\rangle}_{A}\otimes{|{\xi_{A_{1}D}}\rangle}$ results in a zero rate, as $H(A_{1}|D)$ in Proposition 1 becomes non-positive.
@@
The results presented in Theorem 2 and Theorem 3 bear resemblance to the entanglement-assisted quantum capacity of a quantum channel with side information at the transmitter, as studied in Ref. [48]. In the side information model, the channel $\tilde{\mathcal{N}}_{DA\to B}$ is governed by an external environment $D$, which is entangled with the transmitter. The connection becomes apparent if we interpret the transmission to the relay as entangling the transmitter with the channel’s environment (see Fig. 3). Note, however, that the environment state in the side information model [48] is fixed, i.e., $\sigma_{D}$ is dictated by the model. In contrast, our formula involves optimization as we are free to choose the entanglement constellation between the transmitter and the relay.

### v2

@@ IV Main Results
Here we present our main results for the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We establish achievable rates that are based on different coding strategies with and without the entanglement assistance.
@@ IV-A Rate-Limited Entanglement Assistance
Consider the three-party communication scheme in Fig. 3. In the partial decode-forward strategy, the relay does not decode the entire message but only a part of it. Specifically, the quantum “message" register $M$ consists of two subsystems, $M^{\prime}$ and $M^{\prime\prime}$, with the relay decoding $M^{\prime}$ alone. The corresponding transmission rates are denoted by $Q^{\prime}$ and $Q^{\prime\prime}$, so that the overall rate is $Q=Q^{\prime}+Q^{\prime\prime}$.
Before stating the main results, it is important to distinguish between the two tasks for which we establish capacity bounds: quantum information transfer and entanglement generation. As outlined in Sec. I, the full decode-forward strategy addresses quantum information transfer, while the partial decode-forward strategy deals with entanglement generation. The reason for this distinction is structural. In the full decode-forward case, Alice transmits a single quantum message $M$, which is fully decoded by the relay, re-encoded, and forwarded to Bob. Thus, the entire quantum information in $M$ is delivered to the destination receiver. In contrast, in the partial decode-forward case, $M$ is partitioned into two subsystems $M^{\prime}$ and $M^{\prime\prime}$, and only $M^{\prime}$ is decoded and re-encoded at the relay. This partial recovery and forwarding necessitates a final joint decoding at Bob, which reconstructs $M^{\prime}$ and $M^{\prime\prime}$ but not the original undivided system $M$. Consequently, the natural formulation of the partial decode-forward regime is in terms of entanglement generation rather than direct quantum information transfer. We will continue to highlight this distinction in the relevant sections and results that follow.
IV-A1 Achievable Rates
In the proposition below, we assume that there are pre-shared entanglement resources between Alice and Bob, and between Alice and the relay.
@@ IV-A1 Achievable Rates
Consider a quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. Let ${|{\sigma}\rangle}_{AA_{1}A_{2}D}$ be a pure input state, and denote the corresponding output state by $\omega_{A_{1}A_{2}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}A_{2}AD})$. Let $Q^{\prime}\geq 0$ denote the rate of information that is decoded by the relay, and $Q^{\prime\prime}\geq 0$ the rate of information that is sent without the relay’s help. Assume that Alice and Bob have access to entanglement assistance, characterized by the entanglement consumption and generation rates $L^{\prime}_{B},L^{\prime\prime}_{B}\geq 0$ and $\hat{L}^{\prime}_{B},\hat{L}^{\prime\prime}_{B}\geq 0$ in the encoding of $M^{\prime},M^{\prime\prime}$, respectively. Similarly, assume that Alice and the relay have access to entanglement assistance, with consumption and generation rates $L^{\prime}_{E}\geq 0$ and $\hat{L}^{\prime}_{E}\geq 0$. Then any rate point $(Q^{\prime},Q^{\prime\prime},L^{\prime}_{B},\hat{L}^{\prime}_{B},L^{\prime}_{E},\hat{L}^{\prime}_{E},L^{\prime\prime}_{B},\hat{L}^{\prime\prime}_{B})$ such that:
$Q^{\prime}+Q^{\prime\prime}<H(A_{1}A_{2}|D)_{\sigma}-L^{\prime}_{E}-L^{\prime}_{B}-L^{\prime\prime}_{B}$
$\hskip 113.81102pt+\hat{L}^{\prime}_{E}+\hat{L}^{\prime}_{B}+\hat{L}^{\prime\prime}_{B},$
$Q^{\prime}<I(A_{1}\rangle E)_{\omega}+L^{\prime}_{E}-L^{\prime}_{B}-\hat{L}^{\prime}_{E}-\hat{L}^{\prime}_{B},$
$Q^{\prime}+Q^{\prime\prime}<I(A_{1}A_{2}\rangle B)_{\omega}-L^{\prime}_{E}+L^{\prime}_{B}+L^{\prime\prime}_{B}$
$\hskip 113.81102pt-\hat{L}^{\prime}_{E}-\hat{L}^{\prime}_{B}-\hat{L}^{\prime\prime}_{B},$ (9)
is achievable for transmission with rate-limited entanglement assistance, where $A_{1}$ and $A_{2}$ are arbitrary auxiliary systems for the relay and direct links, respectively.
The intepretation of the result above is given in Sec. V, and the proof for Proposition 1 is given in Appendix B.
Intuitively, $A_{1}$ represents Alice’s share of entanglement with the relay, while $A_{2}$ is Alice’s share of entanglement with Bob. Due to entanglement monogamy, the entanglement between $A$ and $A_{1}$ comes at the expense of the entanglement between $A$ and $A_{2}$. As a result, there is a tradeoff between the entanglement that Alice generates with the relay and the entanglement that she generates with the destination decoder.
IV-B Full Decode-Forward Bound (Unassisted)
In the full decode-forward strategy, the relay fully decodes the quantum message received from Alice, re-encodes it, and forwards it to Bob. In that case, the auxiliary register $A_{2}$ is not needed, and the initial state is any pure state ${|{\sigma}\rangle}_{AA_{1}D}$. Our result for the unassisted case is given below. Define
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\min\left\{I(A_{1}\rangle B)_{\omega},I(A_{1}\rangle E)_{\omega}\right\}\right],$ (10)
s.t. $I(A_{1}\rangle B)_{\omega}\geq 0$ and $I(A_{1}\rangle E)_{\omega}\geq 0$, where the maximum is taken over all pure states of the form $|{\sigma_{A_{1}AD}}\rangle$}, and where $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$. The subscript ‘D-F’ stands for decode-forward bound.
@@ IV-B Full Decode-Forward Bound (Unassisted)
$C_{\text{Q}}(\mathcal{N})\geq\mathsf{Q}_{\text{D-F}}(\mathcal{N}).$ (11)
The proof of Theorem 2 is given in Appendix C-A. Intuitively, the formula for in Eq. (10) exhibits a bottleneck behavior, where the worst link dictates the information throughput.
@@
The input state ${|{\sigma}\rangle}_{AA_{1}D}$ represents the correlation that is generated between the transmitter and the relay. Intuitively, we think of the auxiliary system $A_{1}$ as Alice’s entanglement resource. Due to entanglement monogamy, if the resource $A_{1}$ is maximally entangled with the relay’s transmission, $D$, then there is no correlation with the channel input, $A$. As an extreme case, taking ${|{\sigma}\rangle}_{AA_{1}D}={|{\psi}\rangle}_{A}\otimes{|{\phi_{A_{1}D}}\rangle}$ results in a zero rate. Similarly, if the resource is maximally entangled with the channel input, then it has no correlation with the relay. Therefore, it is not immediately obvious how to distribute the entanglement between $_{1}, AA$, and $D$.
@@
Typically, the direct channel from Alice to Bob is noisier than the channel from Alice to the relay, since the destination receiver is located farther away. Consider the case where Bob’s channel is degraded with respect to the relay’s channel. That is, there exists a degrading channel $\mathcal{D}_{E\to B}$ such that $\mathcal{N}_{AD\rightarrow B}=\mathcal{D}_{E\to B}\circ\mathcal{N}_{AD\to E}$. Based on the data processing inequality for the coherent information (see Theorem 11.9.3 in Ref. [63]), the bound reduces to
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})=\max_{\sigma}I(A_{1}\rangle B)_{\omega}.$
@@
IV-C Partial Decode-Forward Bound (Unassisted)
The main result of our work is given below. Define
$\mathsf{Q}_{\text{PD-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\bigl{[}\min\bigl{\{}I(A_{1}A_{2}\rangle B)_{\omega}\,,\\ \;I(A_{1}\rangle E)_{\omega}+I(A_{2}\rangle BA_{1})_{\omega}\bigr{\}}\bigr{]}$ (12)
where the maximum is taken over all pure states of the form $|{\sigma_{A_{1}A_{2}AD}}\rangle$} s.t. $I(A_{1}\rangle B)_{\omega}\geq 0$, $I(A_{1}\rangle E)_{\omega}\geq 0$ and $I(A_{2}\rangle BA_{1})_{\omega}\geq 0$, and where $\omega_{A_{1}A_{2}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}A_{2}AD})$. The subscript ‘PD-F’ stands for partial decode-forward bound.
Theorem 3.
The entanglement-generation capacity of the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$ without assistance satisfies:
$C_{\text{EG}}(\mathcal{N})\geq\mathsf{Q}_{\text{PD-F}}(\mathcal{N}).$ (13)
The proof of Theorem 3 is given in Appendix C-B.
@@ IV-C Partial Decode-Forward Bound (Unassisted)
Theorem 3 can be interpreted as follows. As in Proposition 1, the “message" register is split into two components, $M^{\prime}$ and $M^{\prime\prime}$, at quantum rates $Q^{\prime}$ and $Q^{\prime\prime}$, with a total entanglement-generation rate of $Q=Q^{\prime}+Q^{\prime\prime}$, as defined in Proposition 1. The relay only decodes and encodes the first component, whereas the second component is decoded by Bob without the relay’s help. The rate $Q^{\prime}$ is chosen as
$Q^{\prime}=\min\left\{I(A_{1}\rangle B)_{\omega},I(A_{1}\rangle E)_{\omega}\right\},$
in a similar manner as in the full decode-forward bound (see Eq. (10)), and the rate $Q^{\prime\prime}$ as
$Q^{\prime\prime}=I(A_{2}\rangle BA_{1})_{\omega}.$
The expression $I(A_{1}A_{2}\rangle B)_{\omega}$ in Eq. (12) can be recovered from the entropy chain rule.
Remark 5.
Intuitively, the information rate at the relay $Q^{\prime}$ is determined by the bottleneck, i.e., the weaker link on the relay path. If the Alice-to-relay link is noisier than the relay-to-Bob link, then $I(A_{1}\rangle E)_{\omega}$ limits $Q^{\prime}$. Whereas, if the relay-to-Bob channel is noisier, the limiting term becomes $I(A_{1}\rangle B)_{\omega}$. The formula for $Q^{\prime\prime}$ does not involve the relay output $E$, since the relay is not required to decode this part of the information. While $A_{1}$ cannot be copied due to the no-cloning theorem, it effectively serves as side information at the receiver when decoding $A_{2}$. Indeed, using entropy identities, we can write
$I(A_{2}\rangle BA_{1})_{\omega}=I(A_{2}\rangle B\mid A_{1})_{\omega},$
which can be interpreted as Bob decoding the message component $M^{\prime\prime}$ “given" the message system $M^{\prime}$ sent via the relay.
IV-D Entanglement-Assisted Capacity
In the next theorem, we consider communication of quantum information via the quantum relay channel under the full decode-forward strategy, assuming that Alice and Bob are provided with pre-shared and unlimited entanglement assistance. In this strategy, the relay fully decodes the quantum message, re-encodes it, and forwards it, hence, the auxiliary register $A_{2}$ is not required, and the initial state can be taken as any pure state ${|{\sigma}\rangle}_{AA_{1}D}$ (see also Sec. IV-B). We denote the entanglement-assisted quantum capacity by $C_{\text{Q-EA}}(\mathcal{N})$, where the maximization is as before, and the subscript ‘Q-EA’ indicates the transmission of quantum information using entanglement assistance. This can be viewed as teleportation via a noisy relay channel [64, Sec. 1.3.7].
Let
$\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\frac{1}{2}I(A_{1};B)_{\omega}-\frac{1}{2}I(A_{1};D)_{\sigma}\right].$
Theorem 4.
$C_{\text{Q-EA}}(\mathcal{N})\geq\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N}).$ (14)
The proof of Theorem 4 is given in Appendix C-C. In the analysis, we take $L^{\prime}_{B}>\hat{L}^{\prime}_{B}>0$ and $L^{\prime\prime}_{B}>\hat{L}^{\prime\prime}_{B}>0$ with $L^{\prime}_{E}=\hat{L}^{\prime}_{E}\to 0$ in Proposition 1.
Remark 6.
The theorem above has an immediate consequence on the transmission of classical information with entanglement assistance. Specifically, using the superdense coding protocol [64, Sec. 2.3], classical information can be sent at a rate of $R\geq$ $2\cdot\mathsf{Q}_{\text{EA,PD-F}}(\mathcal{N})$ (classical information bits per transmission).
Remark 7.
@@ IV-D Entanglement-Assisted Capacity
The results bear resemblance to the entanglement-assisted quantum capacity of a quantum channel with side information at the transmitter, as studied in Ref. [51]. In the side information model, the channel $\tilde{\mathcal{N}}_{DA\to B}$ is governed by an external environment $D$, which is entangled with the transmitter. The connection becomes apparent if we interpret the transmission to the relay as entangling the transmitter with the channel’s environment (see Fig. 3). Note, however, that the environment state in the side information model [51] is fixed, i.e., $\sigma_{D}$ is dictated by the model. In contrast, our formula involves optimization as we are free to choose the entanglement constellation between the transmitter and the relay.